# High-Layer Attention Pruning with Rescaling

### v1

@@ 3.3 Long-Context Modeling
In this section, we evaluate the performance of structured pruning algorithms on long-context generation tasks. We use LLaMA3.1-8B and Mistral-8B-v0.3 for our experiments. For evaluation, we use LongBench [7], which encompasses a diverse range of tasks categorized as follows: single-document QA (NrtvQA, Qasper, MF-en), multi-document QA (HotpotQA, 2WikiMQA, Musique), summarization (GovReport, QMSum, MultiNews), few-shot learning (TREC, TriviaQA, SAMSum), synthetic tasks (PCount, PRe), and code-related tasks (Lcc, RB-P). This comprehensive benchmark allows us to assess performance across various long-context scenarios.
@@
Table 3: Performance on long-context generation tasks across LLaMA3.1-8B and Mistral-7B-v0.3.
@@ 3.4 Exploring Layer Selection for Pruning Attention Heads
In this section, we conduct an ablation study to support our claim that attention heads in higher layers are less important than those in lower layers. We evaluate four different strategies for selecting layers to prune, using LLaMA3.1-8B for our experiments. The first strategy prunes attention heads in the lowest 8 layers (Bottom-8). The second selects 8 layers using a Hessian-based metric proposed in [89] (Hessian-8). The third leverages a similarity-based metric for the self-attention module, as proposed in [31], to select 8 layers (Similarity-8). The fourth strategy, which is our approach, prunes attention heads in the highest 8 layers (Top-8). We set $\alpha=1$ for all layers. Additional details on the Hessian-based and similarity-based metrics are provided in Appendix D.
For both the Hessian-8 and Similarity-8 strategies, we compute the respective metrics using the WikiText dataset [63]. The selected layer indices for pruning are [12, 13, 14, 15, 19, 22, 24, 26] based on the Hessian-based metric, and [23, 24, 25, 26, 27, 28, 29, 30] based on the similarity-based metric.
Table 5: Ablation study of HARP on 11 tasks using LLaMA3.1-8B, comparing performance with and without $\alpha$ search (%).
@@ 3.5 $\alpha$ Search for Rescaling
To determine the optimal values of $\alpha$, we compute perplexity on the WikiText [63] and Pile10K [65] datasets using LLaMA3.1-8B. The resulting optimal $\alpha$ values are $[0.8,0.2,0.1,0.1,0.0,0.1,0.0,0.0]$ for WikiText and $[0.8,0.2,0.2,0.1,0.0,0.1,0.1,0.0]$ for Pile10K, with indices corresponding to layers counted from the top. The similarity between the two sets of values suggests a consistent trend in the layer-wise importance.
@@ 3.6 Efficiency Analysis
In this section, we analyze the time complexity of our pruned model. We conduct experiments with varying sequence lengths to measure inference processing time. Using LLaMA3.1-8B as our test model, we evaluate sequences of lengths 1024, 2048, 4096, 8192, 16384, 32768, and 65536. All measurements are conducted on a single NVIDIA H100 80G GPU with each experiment repeated 10 times to ensure statistical reliability. For comparison purposes,
Figure 3: Time complexity analysis across varying sequence lengths, demonstrating efficiency gains of our pruning approach.
we test both our pruned model (with 8 layer attention heads pruned) and the original model (with 0 attention heads pruned) to quantify the efficiency gains. Our measurements reveal consistent performance across all runs, with tight 95% confidence intervals: for the original model, inference times ranged from 0.0371 ± 0.0001s for 1024 tokens to 5.4466 ± 0.0022s for 65536 tokens; for the pruned model, times ranged from 0.0355 ± 0.0001s for 1024 tokens to 4.5394 ± 0.0063s for 65536 tokens. Figure 3 demonstrates the significant speed gains achieved by our pruned model. As sequence length increases, our speed improvement becomes more substantial. With just a 3.3% parameter reduction, we achieve a 16.7% improvement in processing speed for 65,536-token sequences. This highlights how targeted attention head pruning can substantially reduce inference latency in long-context tasks.
@@ 4 Related Work
Pruning is a widely adopted and efficient technique in both Computer Vision and Large Language Models. It can be categorized into two main types: structured pruning and unstructured pruning. structured pruning  [51, 84, 49, 32, 83] involves removing entire filters from neural networks, making it particularly conducive to model deployment. On the other hand, unstructured pruning [13, 70] focuses on removing individual neurons within the network. Some recent works [62, 101, 27, 31, 72, 57, 95, 41, 14, 46] have been proposed to prune blocks in the higher layers of LLMs.
@@
Quantization approaches [82, 91, 23, 85, 19, 68, 55] compress language models by converting weights and activations to lower precision formats such as 8-bit, 4-bit, or even 2-bit integers. This reduction in precision substantially decreases memory requirements and enhances inference speed, enabling the deployment of large language models in environments with limited computational resources.
@@
Recent works [59, 3, 99, 25, 86, 47, 98, 66, 77, 60, 20, 94, 10, 58, 56, 35, 75, 12, 73, 43, 53] have focused on compressing the KV cache to reduce GPU memory footprint. Some [60, 35, 94] of these methods utilize quantization techniques to achieve compression. Additionally, approaches like H2O [99] evict less important tokens, while LESS [20] generates condensed representations through training to optimize memory usage.
@@
Low-rank approximation methods [36, 18, 54, 79, 11] project token representations into a low-dimensional latent space and subsequently reconstruct the full representations via up-projection. This class of methods significantly reduces the number of trainable parameters during fine-tuning.
@@ 5 Conclusion
In this work, we propose HARP, a novel pruning algorithm that removes attention heads in the higher layers of LLMs. Extensive experiments across short-context and long-context generation tasks, as well as discriminative benchmarks, demonstrate the superiority of our proposed method. Ablation studies further reveal the importance of incorporating a rescaling parameter to maintain performance. In addition, efficiency analysis shows that significant gains in long-sequence processing can be achieved by pruning only a small fraction of parameters.
@@ Appendix A Theoretical Analysis
[71] Define :=$\mathcal{M}:=\{\mathbf{Y}\in\mathbb{R}^{N\times d}|\mathbf{Y}=\mathbf{eC},\mathbf{C}\in\mathbb{R}^{1\times d}\}$ as a subspace in $\mathbb{R}^{N\times d}$, where $\mathbf{e}=[1,1,\dots,1]^{\top}\in\mathbb{R}^{N\times 1}$, $N$ is the number of tokens and $d$ is the dimension of token representation.
Clearly, all $\mathbf{Y}$ within the $\mathcal{M}$ space are prone to over-smoothing, as the representations of all tokens in $\mathbf{Y}$ are the same.
@@
[71] Define the distance between matrix $\mathbf{H}\in\mathbb{R}^{N\times d}$ and $\mathcal{M}$ as $d_{\mathcal{M}}(H):=\min_{\mathbf{Y}\in\mathcal{M}}\|\mathbf{H}-\mathbf{Y}\|_{F}$, where $\|\cdot\|_{F}$ is the Frobenius norm.
@@
[71] For self-attention matrix $\hat{\mathbf{A}}$, any $\mathbf{H},\mathbf{B}\in\mathbb{R}^{N\times d}$ and $\alpha_{1},\alpha_{2}\geq 0$, we have:
$d_{\mathcal{M}}(\mathbf{HW})$ $\leq s\,d_{\mathcal{M}}(\mathbf{H}),$ $d_{\mathcal{M}}(\sigma(\mathbf{H}))$ $\leq d_{\mathcal{M}}(\mathbf{H}),$ (6)
$d_{\mathcal{M}}(\alpha_{1}\mathbf{H}+\alpha_{2}\mathbf{B})$ $\leq\alpha_{1}d_{\mathcal{M}}(\mathbf{H})+\alpha_{2}d_{\mathcal{M}}(\mathbf{B}),$ $d_{\mathcal{M}}(\hat{\mathbf{A}}\mathbf{H})$ $\leq\sqrt{\lambda_{\max}}\,d_{\mathcal{M}}(\mathbf{H})$ (7)
By Lemma 1, we have the following lemma.
Lemma 2.
[71] For a Transformer block with $h$ heads, we have
where $v=(1+s^{2})(1+\sqrt{\lambda}hs)/(\beta_{1}\beta_{2})$, $s>0$ is the largest element of all singular values of all $\mathbf{W}^{(\ell)}$, $\lambda$ is the largest eigenvalue of all $\hat{\mathbf{A}}^{\top}(\mathbf{I}-\mathbf{ee}^{\top})\hat{\mathbf{A}}$ for each self-attention matrix $\hat{\mathbf{A}}$, and $\beta_{1}$, $\beta_{2}$ are the minimum standard deviation for two layer normalization operations.
The proofs of these two lemmas can be found in [71]. Lemma 2 shows that if $v<1$, $\mathbf{H}^{(\ell+1)}$ is closer to $\mathcal{M}$ than $\mathbf{H}^{(\ell)}$.
Definition 1. Define the average pairwise cosine similarity among tokens in a matrix $\mathbf{H}\in\mathbb{R}^{N\times d}$ as
$\operatorname{Sim}(\mathbf{H}):=\frac{2}{N(N-1)}\sum_{1\leq i<j\leq N}\frac{\mathbf{h}_{i}^{\top}\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|\|\mathbf{h}_{j}\|}.$ (9)
This measures the global similarity of token representations.
Lemma 3.
For any token representation matrix $\mathbf{H}\in\mathbb{R}^{N\times d}$, the distance to the over-smoothing subspace $\mathcal{M}$ is inversely related to the average pairwise similarity:
@@
By definition, =$\mathcal{M}=\{\mathbf{Y}\in\mathbb{R}^{N\times d}\mid\mathbf{Y}=\mathbf{eC},\mathbf{C}\in\mathbb{R}^{1\times d}\}$, where all token embeddings are equal. If $d_{\mathcal{M}}(\mathbf{H})\to 0$, then there exists $\mathbf{C}$ such that $\mathbf{H}\to\mathbf{eC}$, meaning $\mathbf{h}_{i}\to\mathbf{C}$ for all $i$. Hence, $\mathbf{h}_{i}^{\top}\mathbf{h}_{j}\to\|\mathbf{C}\|^{2}$ and $\|\mathbf{h}_{i}\|\|\mathbf{h}_{j}\|\to\|\mathbf{C}\|^{2}$, so
$\frac{\mathbf{h}_{i}^{\top}\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|\|\mathbf{h}_{j}\|}\to 1,$
and thus $\operatorname{Sim}(\mathbf{H})\to 1$. ∎
Theorem 1. As $\operatorname{Sim}(\mathbf{H}^{(\ell)})\to 1$ for a high-layer $\ell$ in a Transformer, the attention scores in that layer tend to uniform distributions, reducing their discriminative capacity. Therefore, the attention mechanism in higher layers becomes less informative than in lower layers where $\operatorname{Sim}(\mathbf{H}^{(\ell)})<1$.
@@
Suppose $\operatorname{Sim}(\mathbf{H}^{(\ell)})\to 1$, then by Lemma 3, all token embeddings $\mathbf{h}_{i}^{(\ell)}$ converge to a shared direction. Since query and key matrices in self-attention are linear projections of $\mathbf{H}^{(\ell)}$, the inner products $\mathbf{q}_{i}^{\top}\mathbf{k}_{j}$ for all $i,j$ converge to the same value. Therefore, the attention weights
$\text{softmax}(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})$
tend to a uniform distribution, indicating that the attention mechanism can no longer distinguish between different tokens. Consequently, such layers contribute little to feature refinement or discrimination, implying reduced importance. ∎

### v2

@@ 3.3 Long-Context Modeling
In this section, we evaluate the performance of structured pruning algorithms on long-context generation tasks. We use LLaMA3.1-8B and Mistral-8B-v0.3 for our experiments. For evaluation, we use LongBench (Bai et al., 2024), which encompasses a diverse range of tasks categorized as follows: single-document QA (NrtvQA, Qasper, MF-en), multi-document QA (HotpotQA, 2WikiMQA, Musique), summarization (GovReport, QMSum, MultiNews), few-shot learning (TREC, TriviaQA, SAMSum), synthetic tasks (PCount, PRe), and code-related tasks (Lcc, RB-P). This comprehensive benchmark allows us to assess performance across various long-context scenarios.
@@
Table 4: Performance on long-context generation tasks across LLaMA3.1-8B and Mistral-7B-v0.3. Best is bold
@@ 3.4 Exploring Layer Selection for Pruning Attention Heads
In this section, we conduct an ablation study to support our claim that attention heads in higher layers are less important than those in lower layers. We evaluate four different strategies for selecting layers to prune, using LLaMA3.1-8B for our experiments. The first strategy prunes attention heads in the lowest 8 layers (Bottom-8). The second selects 8 layers using a Hessian-based metric proposed in (Yang et al., 2023a) (Hessian-8). The third leverages a similarity-based metric for the self-attention module, as proposed in (He et al., 2024), to select 8 layers (Similarity-8). The fourth strategy, which is our approach, prunes attention heads in the highest 8 layers (Top-8). We set $\alpha=1$ for all layers. Additional details on the Hessian-based and similarity-based metrics are provided in Appendix D.
For both the Hessian-8 and Similarity-8 strategies, we compute the respective metrics using the WikiText dataset (Merity et al., 2017). The selected layer indices for pruning are [12, 13, 14, 15, 19, 22, 24, 26] based on the Hessian-based metric, and [23, 24, 25, 26, 27, 28, 29, 30] based on the similarity-based metric.
Table 6: Ablation study of HARP on 11 tasks using LLaMA3.1-8B, comparing performance with and without $\alpha$ search (%). Best is bold
@@ 3.5 $\alpha$ Search for Rescaling
To determine the optimal values of $\alpha$, we compute perplexity on the WikiText (Merity et al., 2017) and Pile10K (Nanda, 2022) datasets using LLaMA3.1-8B. The resulting optimal $\alpha$ values are $[0.8,0.2,0.1,0.1,0.0,0.1,0.0,0.0]$ for WikiText and $[0.8,0.2,0.2,0.1,0.0,0.1,0.1,0.0]$ for Pile10K, with indices corresponding to layers counted from the top. The similarity between the two sets of values suggests a consistent trend in the layer-wise importance.
@@
Figure 3: Time complexity and runtime scaling with sequence length for the original and pruned models. Solid lines show the absolute runtime (left y-axis), while the green dashed line denotes the relative speedup (%) of the pruned model over the original, plotted on the secondary y-axis (right).
@@ 3.6 Efficiency Analysis
In this section, we analyze the time complexity of our pruned model. We conduct experiments with varying sequence lengths to measure inference processing time. Using LLaMA3.1-8B as our test model, we evaluate sequences of lengths 1024, 2048, 4096, 8192, 16384, 32768, and 65536. All measurements are conducted on a single NVIDIA H100 80G GPU with each experiment repeated 10 times to ensure statistical reliability. For comparison purposes, we test both our pruned model (with 8 layer attention heads pruned) and the original model (with 0 attention heads pruned) to quantify the efficiency gains. Our measurements reveal consistent performance across all runs, with tight 95% confidence intervals: for the original model, inference times ranged from 0.0371 ± 0.0001s for 1024 tokens to 5.4466 ± 0.0022s for 65536 tokens; for the pruned model, times ranged from 0.0355 ± 0.0001s for 1024 tokens to 4.5394 ± 0.0063s for 65536 tokens. Figure 3 demonstrates the significant speed gains achieved by our pruned model. As sequence length increases, our speed improvement becomes more substantial. With just a 3.3% parameter reduction, we achieve a 16.7% improvement in processing speed for 65,536-token sequences. This highlights how targeted attention head pruning can substantially reduce inference latency in long-context tasks.
@@ 4 Related Work
Pruning is a widely adopted and efficient technique in both Computer Vision and Large Language Models. It can be categorized into two main types: structured pruning and unstructured pruning. structured pruning  (Lagunas et al., 2021; Xia et al., 2022; Kurtic et al., 2023; He and Xiao, 2023; Xia et al., 2024) involves removing entire filters from neural networks, making it particularly conducive to model deployment. On the other hand, unstructured pruning (Chen et al., 2020; Sanh et al., 2020) focuses on removing individual neurons within the network. Some recent works (Men et al., 2024; Zhong et al., 2024; Gromov et al., 2024; He et al., 2024; Siddiqui et al., 2024; Liu et al., 2024b; Zhang et al., 2024a; Jaiswal et al., 2024b; Chen et al., 2024b; Kim et al., 2024a) have been proposed to prune blocks in the higher layers of LLMs.
@@
Quantization approaches (Wei et al., 2022b; Yao et al., 2022; Frantar et al., 2023; Xiao et al., 2023; Dettmers and Zettlemoyer, 2023; Park et al., 2024; Lin et al., 2024) compress language models by converting weights and activations to lower precision formats such as 8-bit, 4-bit, or even 2-bit integers. This reduction in precision substantially decreases memory requirements and enhances inference speed, enabling the deployment of large language models in environments with limited computational resources.
@@
Recent works (Liu et al., 2023; Anagnostidis et al., 2023; Zhang et al., 2023; Ge et al., 2024; Xiao et al., 2024; Kim et al., 2024b; Zhang et al., 2024c; Nawrot et al., 2024; Tang et al., 2024; Liu et al., 2024d; Dong et al., 2024a; Yue et al., 2024; Cai et al., 2024; Liu et al., 2024c; a; Hooper et al., 2024; Sun et al., 2024c; Chen et al., 2024a; Sun et al., 2024a; Jiang et al., 2024; Li et al., 2024) have focused on compressing the KV cache to reduce GPU memory footprint. Some (Liu et al., 2024d; Hooper et al., 2024; Yue et al., 2024) of these methods utilize quantization techniques to achieve compression. Additionally, approaches like H2O (Zhang et al., 2023) evict less important tokens, while LESS (Dong et al., 2024a) generates condensed representations through training to optimize memory usage.
@@
Low-rank approximation methods (Hu et al., 2022; Dettmers et al., 2023; Lin et al., 2025; Wang et al., 2025; Chang et al., 2025) project token representations into a low-dimensional latent space and subsequently reconstruct the full representations via up-projection. This class of methods significantly reduces the number of trainable parameters during fine-tuning.
@@ 5 Conclusion
In this work, we propose HARP, a novel pruning algorithm that removes attention heads in the higher layers of LLMs. Extensive experiments across short-context and long-context generation tasks, as well as discriminative benchmarks, demonstrate the superiority of our proposed method. Ablation studies further reveal the importance of incorporating a rescaling parameter to maintain performance.
Beyond these empirical gains, HARP has clear practical implications: by reducing inference cost without requiring architectural changes or retraining from scratch, it can facilitate the deployment of LLMs in resource-constrained environments, enable faster serving for interactive applications, and make long-context reasoning more affordable in real-world systems.
This study has limitations: we mainly focus on attention-head pruning in higher layers, leaving other components underexplored. Future work includes extending HARP to larger and multimodal models, combining it with techniques such as quantization or KV-cache compression, and developing more task-aware pruning criteria to further improve robustness and efficiency.
@@ Appendix A Theoretical Analysis
(Shi et al., 2022) Define the over-smoothing subspace
$\mathcal{M}:=\left\{\mathbf{Y}\in\mathbb{R}^{N\times d}\,\middle|\,\mathbf{Y}=\mathbf{eC},\ \mathbf{C}\in\mathbb{R}^{1\times d}\right\},$
where $\mathbf{e}=[1,1,\dots,1]^{\top}\in\mathbb{R}^{N\times 1}$, $N$ is the number of tokens, and $d$ is the dimension of token representations.
Intuitively, any matrix $\mathbf{Y}\in\mathcal{M}$ has identical rows: every token in the sequence shares exactly the same representation vector. We therefore regard $\mathcal{M}$ as the set of “maximally over-smoothed” token representations, where no token-specific information is preserved.
@@
(Shi et al., 2022) Define the distance between a matrix $\mathbf{H}\in\mathbb{R}^{N\times d}$ and the over-smoothing subspace $\mathcal{M}$ as
$d_{\mathcal{M}}(\mathbf{H}):=\min_{\mathbf{Y}\in\mathcal{M}}\left\|\mathbf{H}-\mathbf{Y}\right\|_{F},$
where $\|\cdot\|_{F}$ denotes the Frobenius norm.
Lemma 1 ((Shi et al., 2022)).
For a self-attention matrix $\hat{\mathbf{A}}$, any $\mathbf{H},\mathbf{B}\in\mathbb{R}^{N\times d}$ and any $\alpha_{1},\alpha_{2}\geq 0$, we have:
$d_{\mathcal{M}}(\mathbf{H}\mathbf{W})$ $\leq s\,d_{\mathcal{M}}(\mathbf{H}),$ $d_{\mathcal{M}}(\sigma(\mathbf{H}))$ $\leq d_{\mathcal{M}}(\mathbf{H}),$ (4)
$d_{\mathcal{M}}(\alpha_{1}\mathbf{H}+\alpha_{2}\mathbf{B})$ $\leq\alpha_{1}d_{\mathcal{M}}(\mathbf{H})+\alpha_{2}d_{\mathcal{M}}(\mathbf{B}),$ $d_{\mathcal{M}}(\hat{\mathbf{A}}\mathbf{H})$ $\leq\sqrt{\lambda_{\max}}\,d_{\mathcal{M}}(\mathbf{H}),$ (5)
By Lemma 1, we have the following result.
Lemma 2 ((Shi et al., 2022)).
For a Transformer block with $h$ heads, we have
where $v=(1+s^{2})(1+\sqrt{\lambda}hs)/(\beta_{1}\beta_{2})$, $s>0$ is the largest element among all singular values of all $\mathbf{W}^{(\ell)}$, $\lambda$ is the largest eigenvalue of all $\hat{\mathbf{A}}^{\top}(\mathbf{I}-\mathbf{ee}^{\top})\hat{\mathbf{A}}$ for each self-attention matrix $\hat{\mathbf{A}}$, and $\beta_{1}$, $\beta_{2}$ are the minimum standard deviations for the two layer normalization operations.
The proofs of these two lemmas can be found in (Shi et al., 2022). Lemma 2 shows that if $v<1$, then $d_{\mathcal{M}}(\mathbf{H}^{(\ell+1)})<d_{\mathcal{M}}(\mathbf{H}^{(\ell)})$, i.e., the hidden representations move closer to $\mathcal{M}$ as the depth $\ell$ increases. In other words, deeper layers become progressively more over-smoothed.
Definition 1. (Average pairwise cosine similarity) Define the average pairwise cosine similarity among tokens in a matrix $\mathbf{H}\in\mathbb{R}^{N\times d}$ as
$\operatorname{Sim}(\mathbf{H}):=\frac{2}{N(N-1)}\sum_{1\leq i<j\leq N}\frac{\mathbf{h}_{i}^{\top}\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|_{2}\|\mathbf{h}_{j}\|_{2}}.$ (7)
A value $\operatorname{Sim}(\mathbf{H})\approx 1$ indicates that all token vectors are nearly parallel (up to scaling), while smaller values correspond to more diverse token directions. Thus, $\operatorname{Sim}(\mathbf{H})$ measures the global similarity of token representations.
@@
For any token representation matrix $\mathbf{H}\in\mathbb{R}^{N\times d}$, if its distance to the over-smoothing subspace $\mathcal{M}$ tends to zero, then its average pairwise similarity tends to one, i.e.,
@@
This is the proof of Lemma 3. By definition, $\mathcal{M}=\{\mathbf{Y}\in\mathbb{R}^{N\times d}\mid\mathbf{Y}=\mathbf{eC},\mathbf{C}\in\mathbb{R}^{1\times d}\}$, where all token embeddings are equal (each row of $\mathbf{Y}$ is $\mathbf{C}$). If $d_{\mathcal{M}}(\mathbf{H})\to 0$, then there exists some $\mathbf{C}$ such that $\mathbf{H}\to\mathbf{eC}$, meaning $\mathbf{h}_{i}\to\mathbf{C}$ for all $i$. Hence,
$\mathbf{h}_{i}^{\top}\mathbf{h}_{j}\to\|\mathbf{C}\|_{2}^{2}\quad\text{and}\quad\|\mathbf{h}_{i}\|_{2}\|\mathbf{h}_{j}\|_{2}\to\|\mathbf{C}\|_{2}^{2},$
so for any $i\neq j$ we have
$\frac{\mathbf{h}_{i}^{\top}\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|_{2}\|\mathbf{h}_{j}\|_{2}}\to 1.$
Averaging over all pairs $(i,j)$ then yields $\operatorname{Sim}(\mathbf{H})\to 1$. ∎
Theorem 1. When $\operatorname{Sim}(\mathbf{H}^{(\ell)})\to 1$, token representations in layer $\ell$ become nearly parallel. Since queries and keys are linear projections of $\mathbf{H}^{(\ell)}$, their dot products across positions become almost constant, so the row-wise softmax yields nearly uniform attention weights. The head thus degenerates into averaging value vectors and cannot distinguish tokens. Conversely, if $\operatorname{Sim}(\mathbf{H}^{(\ell)})$ is bounded away from $1$, attention remains non-uniform and can model meaningful token-to-token dependencies.
@@
This is the proof of Theorem 1. Suppose $\operatorname{Sim}(\mathbf{H}^{(\ell)})\to 1$. By Lemma 3, this implies $d_{\mathcal{M}}(\mathbf{H}^{(\ell)})\to 0$, so all token embeddings $\mathbf{h}_{i}^{(\ell)}$ converge to a shared direction, i.e., they become almost parallel.
In a self-attention layer, the query and key matrices are obtained from $\mathbf{H}^{(\ell)}$ by linear projections:
$\mathbf{Q}^{(\ell)}=\mathbf{H}^{(\ell)}\mathbf{W}_{Q},\qquad\mathbf{K}^{(\ell)}=\mathbf{H}^{(\ell)}\mathbf{W}_{K}.$
Since linear maps preserve the property of vectors being almost parallel, the rows of $\mathbf{Q}^{(\ell)}$ and $\mathbf{K}^{(\ell)}$ also become nearly parallel when $\operatorname{Sim}(\mathbf{H}^{(\ell)})\to 1$. Consequently, the inner products $\mathbf{q}_{i}^{\top}\mathbf{k}_{j}$ for all $i,j$ converge to (almost) the same value:
$\mathbf{q}_{i}^{\top}\mathbf{k}_{j}\approx c\quad\text{for all }i,j.$
The attention weights for a fixed query position $i$ are given by
$a_{ij}=\frac{\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j}/\sqrt{d_{k}})}{\sum_{j^{\prime}}\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j^{\prime}}/\sqrt{d_{k}})}.$
If all logits $\mathbf{q}_{i}^{\top}\mathbf{k}_{j}$ are (approximately) equal, then the softmax output $a_{i1},\dots,a_{iN}]$ approaches the uniform distribution over $\{1,\dots,N\}$. Therefore, the attention weights can no longer distinguish between different tokens, and the corresponding attention head effectively reduces to averaging the value vectors across positions.
As a result, when $\operatorname{Sim}(\mathbf{H}^{(\ell)})$ is close to $1$ in high layers, self-attention behaves like a simple smoothing operator that mixes already similar representations, contributing little to further feature refinement or discrimination. In contrast, in lower layers where $\operatorname{Sim}(\mathbf{H}^{(\ell)})$ is bounded away from $1$, attention weights remain non-uniform and can capture meaningful token-to-token dependencies, making those layers more informative. ∎
Proof of Proposition 1.
Proof.
Recall that $\hat{\mathbf{A}}\in\mathbb{R}^{N\times N}$ is a row-normalized attention matrix and
$\mathbf{H}^{\prime}=\hat{\mathbf{A}}\mathbf{H},\quad\mathbf{H}\in\mathbb{R}^{N\times d}.$
We first show that if every row of $\hat{\mathbf{A}}$ is a one-hot vector (i.e., has exactly one entry equal to $1$ and all others equal to 0), then the Frobenius norm is preserved for any $\mathbf{H}$. In this case, $\hat{\mathbf{A}}$ is a row-permutation matrix: there exists a permutation $\pi$ of $\{1,\dots,N\}$ such that
$(\hat{\mathbf{A}}\mathbf{H})_{i,:}=\mathbf{H}_{\pi(i),:},\quad\forall i.$
Thus $\hat{\mathbf{A}}$ only permutes the rows of $\mathbf{H}$, and we have
$\|\hat{\mathbf{A}}\mathbf{H}\|_{F}^{2}=\sum_{i=1}^{N}\sum_{j=1}^{d}\big((\hat{\mathbf{A}}\mathbf{H})_{ij}\big)^{2}=\sum_{i=1}^{N}\sum_{j=1}^{d}\big(\mathbf{H}_{\pi(i),j}\big)^{2}=\sum_{i=1}^{N}\sum_{j=1}^{d}\mathbf{H}_{ij}^{2}=\|\mathbf{H}\|_{F}^{2}.$
Therefore, if each row of $\hat{\mathbf{A}}$ is one-hot, then $\|\mathbf{H}^{\prime}\|_{F}=\|\mathbf{H}\|_{F}$ for all $\mathbf{H}$.
We now prove the converse: if a row-normalized attention matrix $\hat{\mathbf{A}}$ preserves the Frobenius norm for all inputs, then each row of $\hat{\mathbf{A}}$ must be one-hot. Assume that
$\|\hat{\mathbf{A}}\mathbf{H}\|_{F}=\|\mathbf{H}\|_{F}\quad\text{for all }\mathbf{H}\in\mathbb{R}^{N\times d}.$
Consider the special case where $\mathbf{H}$ has only one nonzero column: let $\mathbf{x}\in\mathbb{R}^{N}$ be arbitrary and define
$\mathbf{H}=\begin{bmatrix}\mathbf{x}&\mathbf{0}&\cdots&\mathbf{0}\end{bmatrix}\in\mathbb{R}^{N\times d}.$
Then
$\hat{\mathbf{A}}\mathbf{H}=\begin{bmatrix}\hat{\mathbf{A}}\mathbf{x}&\mathbf{0}&\cdots&\mathbf{0}\end{bmatrix}$,
and the Frobenius norm condition reduces to
$\|\hat{\mathbf{A}}\mathbf{x}\|_{2}=\|\mathbf{x}\|_{2}\quad\text{for all }\mathbf{x}\in\mathbb{R}^{N}.$
Hence $\hat{\mathbf{A}}$ is an isometry on $\mathbb{R}^{N}$, which implies
$\mathbf{x}^{\top}\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}\mathbf{x}=\|\hat{\mathbf{A}}\mathbf{x}\|_{2}^{2}=\|\mathbf{x}\|_{2}^{2}=\mathbf{x}^{\top}\mathbf{I}\mathbf{x}\quad\forall\mathbf{x},$
so $\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}=\mathbf{I}$. Thus $\hat{\mathbf{A}}$ is an orthogonal matrix.
On the other hand, as an attention matrix, $\hat{\mathbf{A}}$ is row-normalized with nonnegative entries:
$\hat{\mathbf{A}}_{ij}\geq 0,\quad\sum_{j=1}^{N}\hat{\mathbf{A}}_{ij}=1\quad\text{for each row }i.$
Let $\mathbf{a}_{i}$ denote the $i$-th row of $\hat{\mathbf{A}}$. Orthogonality implies $\|\mathbf{a}_{i}\|_{2}=1$ for all $i$, while row-normalization gives $\sum_{j}\mathbf{a}_{ij}=1$ and $\mathbf{a}_{ij}\geq 0$. Therefore,
$\sum_{j=1}^{N}\mathbf{a}_{ij}=1,\quad\sum_{j=1}^{N}\mathbf{a}_{ij}^{2}=1,\quad\mathbf{a}_{ij}\geq 0.$
For any $0<a<1$, we have $a^{2}<a$. Hence, if any entry of $\mathbf{a}_{i}$ satisfies $0<\mathbf{a}_{ij}<1$, then
$\sum_{j=1}^{N}\mathbf{a}_{ij}^{2}<\sum_{j=1}^{N}\mathbf{a}_{ij}=1,$
which contradicts $\sum_{j}\mathbf{a}_{ij}^{2}=1$. Thus every entry of $\mathbf{a}_{i}$ must belong to $\{0,1\}$. Since the entries are nonnegative and the row sum is $1$, each row has exactly one entry equal to $1$ and all others equal to 0, i.e., each row is one-hot.
Combining both directions, we conclude that a row-normalized attention matrix preserves the Frobenius norm of all input representations if and only if each of its rows is a one-hot vector. In practical self-attention, the learned attention weights are not row-permutation matrices, so the aggregation step $\mathbf{H}^{\prime}=\hat{\mathbf{A}}\mathbf{H}$ typically changes the magnitude (Frobenius norm) of the token representations, as claimed in Proposition 1. ∎